OpenStudy (anonymous):

Find the measure of an angle whose supplement is three times its complement.

OpenStudy (anonymous):

let the angle be x

OpenStudy (anonymous):

complement = 90- x supplement = 180-x

OpenStudy (anonymous):

180-x = 3(90-x) 180 - x = 270 - 3x x = 45

OpenStudy (anonymous):

Did u Understand..!!

OpenStudy (anonymous):

Thanks ^^

OpenStudy (anonymous):

Welcome...))

OpenStudy (anonymous):

if u r satisfied plzz give medals